The length, width, and height of a rectangular solid are in ratio of 4:3:2. If the volume of the box is 648, what is the total surface area of the box?
Answer by rfer(16322)

4x*3x*2x=648
24x^3=648
x^3=27
x=3
------------------
4x=12
3x=9
2x=6
-------------------
here are the dimensions, can you figure the total area?